# Contrastive Learning Augmented Social Recommendations

### v2

@@ 3.1. Performance Comparison
Table 1 compares the proposed method with state-of-the-art baselines on the LastFM and Ciao datasets using Precision, Recall, and NDCG metrics. On the LastFM dataset, the proposed method consistently outperforms SocialLGN in most metrics, achieving notable improvements in Precision@10, Recall@10, and NDCG@10, with gains of up to 1.92%. On the Ciao dataset, the proposed method demonstrates significant improvements in Precision and NDCG, particularly at top-ranked positions, with gains of up to 2.81%.
The results highlight the effectiveness of the proposed method in improving top-ranked recommendations. On the LastFM dataset, the method excels in capturing user preferences, particularly in ranking relevant items higher in shorter recommendation lists. On the Ciao dataset, the method demonstrates robustness by achieving substantial gains in Precision and NDCG, which are critical for evaluating top-ranked recommendations. In summary, the proposed method consistently improves top-ranked recommendations, particularly on Precision@10, Recall@10, and NDCG@10, demonstrating its effectiveness in refining social information.
3.2. Cold Users
Table 1 compares methods for addressing the cold-user problem, where users with fewer than 20 interactions are defined as cold users (liao2022sociallgn, ).
On LastFM, CLSRec exhibits significant and substantial relative-gains across all metrics, with improvements exceeding 18% in Precision, 13% in Recall, and 10% in NDCG. On Ciao, the performance gains are more modest, with slight improvements in Recall and NDCG, while Precision shows mixed results depending on the recommendation list length. Overall, CLSRec effectively mitigates the cold-user problem, achieving substantial gains on LastFM and maintaining competitive performance on Ciao, validating its robustness across datasets.
@@ 4. Conclusion
In this paper, we proposed a CLSRec recommendation method to fuse social relation into behavior-based recommendations, which leverages the reconstructed social view to enhance the user interest representation, addressing key challenges like social relation noise and interest inconsistency. By introducing a dual-view augmentation approach, the method effectively filters the social information, mitigating noise and sparsity issues. Additionally, the mutual distillation method aligns and fuses social and behavioral interest representations, ensuring that knowledge from both sources is effectively combined without loss. Experimental results validate the effectiveness of the approach and dramatically improve recommendations for cold users, demonstrating its great potential to enhance user experiences and drive engagement on modern content platforms.

### v3

@@ 3.1. Performance Comparison
Table 1 compares the proposed method with state-of-the-art baselines on the LastFM and Ciao datasets using Precision, Recall, and NDCG metrics. On the LastFM dataset, the proposed method consistently outperforms GBSR in most metrics, achieving notable improvements in Precision@10, Recall@10, and NDCG@10, with gains of up to 1.71%. On the Ciao dataset, the proposed method demonstrates significant improvements in Precision and NDCG, particularly at top-ranked positions, with gains of up to 3.41%.
Additionally, for the cold-user, where users with fewer than 20 interactions (liao2022sociallgn, ), CLSRec exhibits significant and substantial relative-gains across all metrics, with improvements exceeding 15% in Precision, 13% in Recall, and 7% in NDCG on LastFM dataset. On Ciao, the performance gains are more modest, with slight improvements in Recall and NDCG, while Precision shows mixed results depending on the recommendation list length. Overall, CLSRec effectively improves top-ranked recommendations and mitigates the cold-user problem, achieving substantial gains on LastFM and Ciao, validating its robustness across datasets.
3.2. Ablation Study
Table 2 shows the ablation study of CLSRec on the LastFM dataset. Removing the contrastive loss (CL) slightly degrades performance, indicating its role in enhancing representation learning. Further removing the interest inconsistency alignment (IIA) leads to a more noticeable drop, demonstrating its importance in capturing complementary information. Overall, both components contribute to the effectiveness of the proposed CLSRec.
Table 2. Ablation Study of CLSRec on LastFM dataset
Method Precision@20 Recall@20 NDCG@20
CLSRec 0.1384 0.2829 0.2872
w/o CL 0.1368 0.2794 0.2883
w/o CL & IIA 0.1358 0.2769 0.2788
@@ 4. Conclusion
In this paper, we proposed a CLSRec recommendation method to fuse social relation into behavior-based recommendations, which leverages the reconstructed social view to enhance the user interest representation, addressing key challenges like social relation noise and interest inconsistency. Experiments validate the effectiveness of the approach and dramatically improve recommendations for cold users. The dual-view modeling and social-interest alignment method establishes a promising paradigm, suggesting potential for broad adoption in social-behavior recommendation for future research.
Acknowledgements.
This project has been supported by the Hong Kong Research Grants Council under General Research Fund (project no. 15200023).